OpenStudy (thadds2003):

What is the solution to the equation? 7(5 + 9) = 7 • x + 7 • 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (gabbyalicorn):

The answer is B. 7(5+9)=7x+(7)(9) Simplify both sides of the equation. 98=7x+63 Flip the equation. 7x+63=98 Subtract 63 from both sides. 7x+63−63=98−63 7x=35 Divide both sides by 7. 7x7=357 Answer: x=5

OpenStudy (thadds2003):

thx, can you help me with 2 more?

OpenStudy (gabbyalicorn):

Sure, as long as it's on algebraic equations, I'll be happy to help you! ^-^

OpenStudy (thadds2003):

ok, here the other q.

OpenStudy (gabbyalicorn):

I made the blank spot x if that's okay for your assignment. Simplify 8×63 to 504 504=8x+8×3 Simplify 8×3 to 24 504=8x+24 Subtract 24 from both sides 504−24=8x Simplify 504−24 to 480 480=8x Divide both sides by 8 4808=x Simplify 4808 to 60 60=x Switch sides x=60 Answer is 60. Double checking: Let x=60 8×63=8×60+8×3 Simplify 8×63 to 504 504=8×60+8×3 Simplify 8×60 to 480 504=480+8×3 Simplify 8×3 to 24 504=480+24 Simplify 480+24 to 504 504=504

OpenStudy (thadds2003):

whats the first blank?

OpenStudy (gabbyalicorn):

The first blank was 60 but let me recalculate.

OpenStudy (gabbyalicorn):

Last blank is 504

OpenStudy (thadds2003):

ok, here last q.

OpenStudy (gabbyalicorn):

Is this a test?

OpenStudy (thadds2003):

no, beyond the lesson, i swear.

OpenStudy (gabbyalicorn):

Hmm, okay.

OpenStudy (gabbyalicorn):

Sorry I took so long. it is 12.2 and 44.0.

OpenStudy (thadds2003):

thx sooooooo much!

OpenStudy (gabbyalicorn):

No problem! ^-^